OpenStudy (anonymous):

Find the equation of the straight line that is parallel to the lines y=6 and y=2 and the midway between them.

OpenStudy (anonymous):

Parallel means same slope. These lines are horizontal, so the slope is 0. It will be in the form of y = k where k is some constant like 6 and 2. Halfway between 6 and 2? That's just the average, so: (6 + 2)/2 which = 8/2 = 4. y = 4

OpenStudy (anonymous):

right, y = 4.

OpenStudy (anonymous):

 y=4